Zelynross:

What is the conjugate of the following complex number: 6 - 2i

Zelynross:

@Ultrilliam

KjSaif:

Ok the first step is that you rewrite (6 - 2i) into


> \[2\times 3 - 2i\]

KjSaif:

next you factor out the common term (2 for this question) \[2(3-i)\]

KjSaif:

If you rewrite this in standard complex form you would get the same answer as orignal, are there any answer choices?

Zelynross:

no

KjSaif:

ok since you can't factor \[6 - 2i\] any further, you instead flip the subtraction with the opposite term, addition.

KjSaif:

This gives you an answer of \[6+2i\]

KjSaif:

All you do is negate the term with "i"

KjSaif:

hope this helps.

Zelynross:

preciate it

KjSaif:

np

Zelynross:

Find the quotient (using complex conjugates) 1 over 6-2i

jhonyy9:

have you understood it what mean conjugate ? \[\frac{ 1 }{ 6 -2i } = \] - so when there is a term with ,,i" in denominator we need eliminate it - so in this way we need multiplie top and bottom by conjugate of denominator how you think what maybe the conjugate of 6 -2i ? use these above wrote @Laylalyssa any idea here pls. ? do you know the way ?

jhonyy9:

@supie any idea here pls ?

Laylalyssa:

nah me no know the way

jhonyy9:

have you learn about complex numbers ?

Laylalyssa:

uhh no..